OpenStudy (anonymous):

a) a box is 5 ft * 3 ft * 2 ft is to be filled with cubes that are 2in * 2in.Find the number of cubes it will take to fill this show your work and explain answers pleasee b) if you reduced each dimension of the box by 25% what would the volume of the resulting box be? show your workk pleasee c) you want to design a box with a volume of 125 cubic ft. give one possible dimensions for your box if all dimensions are integers?

OpenStudy (anonymous):

1 foot = 12 inch a) we have to change the size to inch 60*36*24 inch  at side of 24 we can fit 24/2 = 12 cube at side of 36 we can fit 36 /2= 18 cube at side of 60 which is the elevation 60 / 2 = 30 then number of cubs = 12 * 18 * 30 = 6480 cube b) 5 ft * 3 ft * 2 ft it is reduced by 25% 60 inch * 36 Inch * 24 Inch 60 *0.25 = 15 inch this value is not the new one .. its only mean that the origin tall must reduce by it so 60 - 15 = 45 inch 36 * 0.25 = 9 inch 36 - 9 = 27 inch 24 * 0.25= 6 inch 24 - 6 = 18 inch so the new dim. = 45 inch * 27 inch * 18 Inch or 3.75 ft *2.25 ft *1.5 ft d) 125 cubic ft 5 ft * 5 ft * 5 ft Hope I'm right ... You are welcome ... regards